OpenStudy (anonymous):

When the price of a product is p dollars, each, suppose that a manufacturer will supply 2 p −10 units of the product to the market and that consumers will demand to buy 200 −3p units. At the value of p for which supply equals demand, the market is said to in equilibrium. Find the value of p . What is your reply?

OpenStudy (anonymous):

2p - 10 = 200 - 3p First you will subtract add 2p to both sides. +3p + 3p ---------------------- 5p - 10 = 200 Then i will add 10 to both sides +10 + 10 ---------------- 5p = 210 Now divide both sides by 5 to get the value of p. /5 /5 --------- p = 42

OpenStudy (anonymous):

Do you understand now @s1421?

OpenStudy (anonymous):

Yes thanks a lot :)